OpenStudy (anonymous):

what is the indefinite integral of 4sec4x .tan4x.sec^4 4x dx; u = sec4x and the other qoustion is integral of (e^4x -4)^1/5 .8e^4x .dx plz help me

ganeshie8 (ganeshie8):

ya that wil work for first question : u = sec4x

OpenStudy (anonymous):

YES U=SEC4X

OpenStudy (anonymous):

\[\int8e^{4x}\left(e^{4x}-4\right)^{1/5}~dx\] Substitute \(u=e^{4x}-4\), then \(\dfrac{1}{4}du=e^{4x}~dx\): \[2\int u^{1/5}~du\]